Any continuous function f defined on a closed interval I=[a,b] is uniformly continuous.Prove
1
Expert's answer
2020-10-26T20:21:06-0400

If "f" is not continuous then there exist "\epsilon >0" such that for any "\delta >0," must exist points "x_k, y_k" such that "|x_k-y_k|<\frac{\delta}{2^k}" but "|f(x_k)-f(y_k)|>\epsilon" for each "k>0." Now the sequence is bounded by the interval, hence by Bolzano-Weirstrass a subsequence of "{x_k}" which converges. For simplicity let us call this sequence also as "{x_k}." Now let "x_k\rightarrow x." Then we claim "y_k\rightarrow x." Let "\epsilon'>0" be given. Then there exists "k_1, N>0," such that "\frac{\delta}{2^{k_1}}<\frac{\epsilon^{'}}{2}" and "|x_{k}-x|<\epsilon' \ \forall \ k\geq N." We take "k=max\{k_{1},N\}."

By triangle inequality "|y_k-x|\leq |y_k-x_k|+|x_k-x|""<\frac{\delta}{2^{k_1}}+\frac{\epsilon^{'}}{2} \leq \frac{\delta}{2^k}+\frac{\epsilon'}{2}< \frac{\epsilon'}{2}+\frac{\epsilon'}{2}= \epsilon'."

Now "f" being continuous, "f(y_k)\rightarrow f(x)." Also, "f(x_k)\rightarrow f(x)." So given "\epsilon>0" there exist "N_1,N_2" such that for all "k>N_1,N_2" ",|f(x_k)-f(x)|<\frac{\epsilon}{2}, |f(y_k)-f(x)|<\frac{\epsilon}{2}." Hence "|f(x_k)-f(y_k)|\leq |f(x_k)-f(x)|+|f(y_k)-f(x)|" "\leq \epsilon" contradicting "|f(x_k)-f(y_k)|>\epsilon ."